OpenStudy (anonymous):

What is the y-value of sin(x) when x= -180°?

jimthompson5910 (jim_thompson5910):

sin(x) = sin(-180) ... replace x with -180 sin(-180) = -sin(180) ... use the rule sin(-x) = -sin(x) -sin(180) = -1*sin(180) = -1*0 = 0 I'm using the unit circle on the last step to evaluate sin(180)

OpenStudy (anonymous):

So the answer would be 0?

jimthompson5910 (jim_thompson5910):

correct